OpenStudy (anonymous):

Find the area of a rectangle tennis court if the perimeter is 480 feet. The length of the court is six feet more than twice the width. A. Not given B. A= 12,606 ft 2 C. A= 10,636 ft 2 D. A= 12,636 ft 2

OpenStudy (johnweldon1993):

so the perimeter is 480....lenth is 2w + 6 right?

OpenStudy (anonymous):

area = length *breadth

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

what does breadth mean?

OpenStudy (anonymous):

Yes @johnweldon1993

OpenStudy (johnweldon1993):

lol i believe he meant width

OpenStudy (anonymous):

thats what i thought :)

OpenStudy (ghazi):

2((2w+6)+w)=480

OpenStudy (anonymous):

Thank You ghazi

OpenStudy (johnweldon1993):

yes...as @ghazi as written

OpenStudy (anonymous):

i dont know the length how am i supposed to do anything with this stupid problem :(

OpenStudy (ghazi):

2(3w+6)=480 divide both sides by 2 3w+6 = 240 subtract 6 on both sides 3w=240-6 3w=234 divide by three on both sides w= 78 and your length = 2*78+6=156+6=162

OpenStudy (ghazi):

and now you can have area = 162* 78 feet squared

OpenStudy (anonymous):

:)

OpenStudy (ghazi):

hahahaha

OpenStudy (anonymous):

so @lovemenplayhard area is length times width right?

OpenStudy (anonymous):

yep (; whole solution is right there lol

OpenStudy (ghazi):

hahahahaha

OpenStudy (anonymous):

thanks!! :D so nice of you

OpenStudy (ghazi):

you are welcome :)

OpenStudy (anonymous):

both of you

OpenStudy (anonymous):

anytym :)

OpenStudy (anonymous):

;)

OpenStudy (anonymous):

;) ;) lets tak ;)

OpenStudy (anonymous):

I don't like you I love Ghaziii!!

OpenStudy (ghazi):

i love you too honey ;)

OpenStudy (anonymous):

dats foking rude i dint say i like u i was interstd in tok

OpenStudy (ghazi):

mind your tongue dude :D

OpenStudy (anonymous):

else wat? lol

OpenStudy (ghazi):

alright let see else what :D

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

guys, please...

OpenStudy (anonymous):

WTF u guys r crazy, i got foking warning , wat for ?

OpenStudy (ghazi):

do you want another one lol

OpenStudy (anonymous):

hell no F off @ghazi

OpenStudy (ghazi):

thanks :)

OpenStudy (anonymous):

who did that?? i didnt do anything

OpenStudy (ghazi):

i did :) for bad tongue :D he called your statement with bad words :(

OpenStudy (anonymous):

wateva ayam sorry ayam outta here